# From Static to Adaptive Defense: Federated Multi-Agent Deep Reinforcement Learning-Driven Moving Target Defense Against DoS Attacks in UAV Swarm Networks

### v1

@@ IV-C Federated Parameter Aggregation
After a number of local updates (e.g., $K$ episodes), all agents upload their local parameters $\theta_{i}$ to the central aggregator. Instead of naive averaging, we perform a reward-weighted aggregation of policy parameters. For each participating agent $i$, we compute a weight $w_{i}$ proportional to its recent average return over the most recent $M$ episodes, denoted as:
@@
where $r_{t,m}^{i}$ is the reward received by agent $i$ at time step $t$ in the $m$-th most recent episode. The weight for agent $i$ is then calculated as:
@@
To balance generalization and personalization, we only aggregate the parameters of the shared layers (e.g., feature extraction layers) across agents. The output layers remain local to each agent, allowing for adaptation to individual UAV roles or local environments. Thus, the global parameter for each shared layer is computed as:
@@ V-A Simulation Setup
We investigate a coverage area of $1\times 1$ Km serviced by a single GCS and multiple UAVs. The GCS is located at the center of the area, specifically at coordinates [500, 500, 0] m. Each UAV in our method is responsible for patrolling and monitoring the ground area, and all of them build a formation with a radius of 300 m and a height of 100 m. The minimum separation distance between UAVs is set to 20 m, and the deviation threshold for maintaining formation is set to 40 m. The attack duration is set to 15 s, and the reconnaissance period is set to 5 s. Each MTD mechanism has an execution time of 1 s. Other simulation parameters and their default values are summarized in Table I with reference to [33, 34, 35]. All the experiments were conducted on a workstation with 2.20 GHz Intel Xeon Gold 5220R, NVIDIA RTX A5000, 128 GB RAM, and Ubuntu 20.04 operating system. During the training phase, we trained the models for 2000 episodes and repeated the training process with random seeds ranging from 1 to 10 to ensure statistical reliability. In the testing phase, each method was evaluated over 10 episodes using random seeds from 1 to 100 to reduce the uncertainty of the environment.
@@ V-B Convergence Analysis
TABLE II: Attack Mitigation Rate under Different Attack Types and Strategies
@@ V-C Effectiveness and Efficiency Analysis
Notably, PG-FMADRL performs exceptionally well even against greedy attackers. In the node-level scenario, it achieves a attack mitigation rate of 0.9996, slightly below DESOLATER but 34.6% higher than RE-MTD. In the more challenging link-level scenario, PG-FMADRL reaches an attack mitigation rate of 0.9367, significantly outperforming the other methods, which achieve only 0.7910, 0.7625, 0.7360, and 0.7721, respectively. This result indicates that, compared to other methods, PG-FMADRL is more capable of continuously and adaptively responding to dynamic attackers, enabling more intelligent and effective defense decisions.
@@ V-D Energy Consumption and Defense Cost Analysis
To estimate the energy consumption of the UAV swarm when equipped with different defense methods, we adopt a simplified model introduced by Liu et al. [39] to express the power of a UAV according to its velocity as:
@@
where the coefficients are set to $c_{1}=2.8037\,(\mathrm{m}/\mathrm{kg})^{1/2}$, $c_{2}=0.3177\,(\mathrm{m}/\mathrm{kg})^{1/2}$, $c_{3}=0.0296\,\mathrm{kg}/\mathrm{m}$, the mass of an UAV is set to $m=1.283\,\mathrm{kg}$, and the gravitational acceleration is set to $g=9.8\,\mathrm{m}/\mathrm{s}^{2}$ with reference to [39].
Therefore, the total energy consumption in this section is computed as the sum of the energy consumed during the patrolling and return phases for all UAVs, as illustrated in Fig. 5. From an overall perspective, the greedy attack strategy leads to significantly higher energy consumption for the UAV swarm compared to the fixed and random strategies. This is primarily because the greedy attacker exhibits greater attack intensity and adaptability, dynamically adjusting its targets in response to defense actions. As a result, the swarm is often forced to deviate from its optimal formation, causing UAVs to travel longer distances and thus to consume more energy. Nevertheless, our proposed method consistently achieves the lowest energy consumption across all scenarios. Notably, under node attack conditions, PG-FMADRL reduces energy consumption by 29.3% compared to the WF-MTD method when facing greedy attackers, thereby effectively extending the patrol duration of the UAV swarm.
We can also observe that the energy consumption of DESOLATER is very close to our method under node attacks, which can be attributed to the fact that DESOLATER also employs a policy gradient-based approach. However, in the link attack scenario, PG-FMADRL enables more effective global agent coordination and learns superior action sequences by leveraging the federated learning framework. Consequently, under greedy attack strategies, PG-FMADRL achieves a 10.9% reduction in energy consumption compared to DESOLATER, further highlighting its effectiveness in collaborative defense and energy efficiency.
@@ VI Conclusion and Future Work
In this paper, we proposed a novel FMADRL-driven MTD framework to proactively and adaptively mitigate DoS attacks in UAV swarm networks. By designing three lightweight and coordinated MTD mechanisms (e.g., leader switching, route mutation, and frequency hopping) and formulating the defense problem as a multi-agent POMDP, our approach enables each UAV to autonomously select defense actions based on local observations while benefiting from collaborative learning through reward-weighted federated aggregation. Extensive simulation results demonstrated that the proposed PG-FMADRL method significantly outperforms state-of-the-art baselines in terms of attack mitigation rate, recovery time, energy consumption, and defense cost, while maintaining robust mission continuity under various attack scenarios. Our source code can be available at https://github.com/SEU-ProactiveSecurity-Group/PG-FMADRL.
Despite these promising results, several avenues remain for future research. First, this work primarily considers attackers with fixed, random, or greedy strategies. In practical scenarios, adversaries may also leverage advanced techniques such as DRL to dynamically adapt their attack strategies in response to defense mechanisms. Therefore, a promising direction is to investigate adaptive defense strategies against intelligent attackers that can also employ DRL-based decision-making, leading to a more realistic and challenging adversarial environment. Additionally, future work will explore the scalability of the proposed framework to larger-scale UAV swarms and more complex mission scenarios, as well as the integration of additional lightweight MTD mechanisms to further enhance system robustness and security.

### v2

@@ IV-C Federated Parameter Aggregation
After a number of local updates (e.g., $K$ episodes), all agents upload their local parameters $\theta_{i}$ to the central aggregator. Instead of naive averaging, we perform a reward-weighted aggregation of policy parameters. For each participating agent $i$, we compute a weight $w_{i}$ proportional to its recent average return over the most recent $M$ episodes, denoted as
@@
where $r_{t,m}^{i}$ is the reward received by agent $i$ at time step $t$ in the $m$-th most recent episode. The weight for agent $i$ is then calculated as
@@
To balance generalization and personalization, we only aggregate the parameters of the shared layers (e.g., initial hidden layers) while the final output layer is kept local. The boundary between them is determined by the principle of separating general knowledge from specialized decision-making, allowing each agent to personalize its final action-selection mapping based on its specific role and recent experiences. Thus, the global parameter for each shared layer is computed as
@@ V Performance Evaluation
This section evaluates the effectiveness and efficiency of our proposed PG-FMADRL method through extensive simulations. We first describe the simulation setup and baseline methods used for comparison. Then, we analyze the convergence properties of our algorithm, followed by a comprehensive assessment of its performance in terms of attack mitigation rate, recovery time, energy consumption, and defense cost under various attack scenarios. Finally, we assess the scalability of our approach compared to other SOTA methods.
@@ V-A Simulation Setup
We investigate a coverage area of $1\times 1$ Km serviced by a single GCS and multiple UAVs. The GCS is located at the center of the area, specifically at coordinates [500, 500, 0] m. Each UAV in our method is responsible for patrolling and monitoring the ground area, and all of them build a formation with a radius of 300 m and a height of 100 m. For our primary experiments (Sections V-B to V-D), we use a configuration with $N=5$ UAVs for a detailed and fair comparison across multiple metrics and baselines. To specifically address scalability, we then conduct an analysis in Section V-E with varying swarm sizes. The minimum separation distance between UAVs is set to 20 m, and the deviation threshold for maintaining formation is set to 40 m. The attack duration is set to 15 s, and the reconnaissance period is set to 5 s. Each MTD mechanism has an execution time of 1 s.
The key hyperparameters, including the reward weighting coefficients $(\alpha,\beta,\zeta,\eta,\xi)$ and the entropy coefficient $\mu$, were determined through an iterative, empirical tuning process. This process aimed to find a balance between competing objectives, such as prioritizing attack mitigation over defense cost and managing the exploration-exploitation trade-off, to ensure the policy converges to a stable and effective defense strategy. The final values for all simulation parameters are summarized in Table I with reference to [38, 39, 40].
All the experiments were conducted on a workstation with 2.20 GHz Intel Xeon Gold 5220R, NVIDIA RTX A5000, 128 GB RAM, and Ubuntu 20.04 operating system. During the training phase, we trained the models for 2000 episodes and repeated the training process with random seeds ranging from 1 to 10 to ensure statistical reliability. In the testing phase, each method was evaluated over 10 episodes using random seeds from 1 to 100 to reduce the uncertainty of the environment.
@@
It is worth noting that, although some of the baseline methods were originally proposed for general network environments or cloud networks rather than UAV-specific scenarios, we have carefully adapted and implemented all baseline methods within the UAV swarm context. Specifically, each baseline is configured to operate under the same UAV swarm system model as our proposed approach, and all utilize the same set of lightweight MTD mechanisms as described in Section III-C. The original algorithms and decision-making frameworks of each baseline are retained, and all methods are trained and evaluated using the same simulation environment and attack scenarios to ensure consistency and fairness in the comparative evaluation.
@@ V-B Convergence Analysis
In the end, the rapid and stable convergence of PG-FMADRL can be attributed to its federated learning framework, which enables UAV agents to share policy parameters and aggregate knowledge from diverse local experiences. This collaborative approach accelerates learning, helps avoid local optima, and ensures that the swarm quickly adapts to various attack strategies, especially in dynamic and partially observable environments.
TABLE II: Attack Mitigation Rate under Different Attack Types and Strategies. The highest average mitigation rates (green) and the lowest standard deviations (orange) for each attack are highlighted.
@@ V-C Effectiveness and Efficiency Analysis
Notably, PG-FMADRL’s advantage becomes most apparent against sophisticated greedy attackers. While ID-HAM and DESOLATER rely on independent learning where each agent must discover effective strategies in isolation, our federated architecture acts as a knowledge accelerator. For instance, in the challenging greedy link attack scenario, PG-FMADRL reaches an attack mitigation rate of 0.9367, significantly outperforming the other methods, which achieve only 0.7910, 0.7625, 0.7360, and 0.7721, respectively. This is because our reward-weighted aggregation mechanism allows the entire swarm to rapidly learn from the experiences of the most successful agents those achieving high rewards, by effectively countering the attack. Our collaborative learning ensures that effective defense policies are quickly propagated and adopted across the swarm, leading to a more robust and coordinated response.
@@
In summary, the superior attack mitigation rates and reduced recovery times of PG-FMADRL stem from its reward-weighted aggregation and distributed policy learning. This enables UAVs to collaboratively reinforce successful defense behaviors and adaptively coordinate MTD actions in real time. Compared to other multi-agent methods such as DESOLATER and ID-HAM, this approach ensures that high-performing agents have greater influence on the global policy, leading to more robust and consistent defense against dynamic and adaptive attacks.
(a) Node attack
(b) Link attack
@@ V-D Energy Consumption and Defense Cost Analysis
To estimate the energy consumption of the UAV swarm when equipped with different defense methods, we adopt a simplified model introduced by Liu et al. [44] to express the power of a UAV according to its velocity, which includes three components, namely induced power, profile power, and parasite power as
@@
where $c_{1}$ and $c_{2}$ represent empirical coefficients related to induced and profile power ($c_{1}=2.8037\,(\mathrm{m}/\mathrm{kg})^{1/2}$, $c_{2}=0.3177\,(\mathrm{m}/\mathrm{kg})^{1/2}$), $c_{3}$ is the parasitic power coefficient ($c_{3}=0.0296\,\mathrm{kg}/\mathrm{m}$), $m$ is the mass of the UAV ($m=1.283\,\mathrm{kg}$), and $g$ is the gravitational acceleration ($g=9.8\,\mathrm{m}/\mathrm{s}^{2}$) with reference to [44].
This model allows us to quantitatively compute the consumed power and compare the energy efficiency of each defense strategy, and the results are presented in Fig. 5. From an overall perspective, the greedy attack strategy leads to significantly higher energy consumption for the UAV swarm compared to the fixed and random strategies. This is primarily because the greedy attacker exhibits greater attack intensity and adaptability, dynamically adjusting its targets in response to defense actions. As a result, the swarm is often forced to deviate from its optimal formation, causing UAVs to travel longer distances and thus to consume more energy. Nevertheless, PG-FMADRL consistently achieves the lowest energy consumption across all scenarios. Notably, under node attack conditions, it reduces energy consumption by 29.3% compared to WF-MTD when facing greedy attackers, thereby effectively extending the patrol duration of the swarm.
We can also observe that the energy consumption of DESOLATER is very close to our method under node attacks, which can be attributed to the fact that DESOLATER also employs a policy gradient-based approach. However, in the link attack scenario, PG-FMADRL enables more effective agent coordination by leveraging the federated learning framework. Consequently, under greedy attack strategies, PG-FMADRL achieves a 10.9% reduction in energy consumption compared to DESOLATER, further highlighting its effectiveness in collaborative defense and energy efficiency.
@@
The lower energy consumption and defense cost achieved by PG-FMADRL are attributed to its intelligent and context-aware action selection. By leveraging policy gradient optimization, each UAV can precisely determine when and which MTD mechanism to trigger, thus minimizing unnecessary or redundant actions. This fine-grained control ensures efficient use of resources while maintaining strong protection, outperforming baseline methods that often rely on more frequent or less targeted defense operations.
V-E Scalability Analysis
A critical aspect of any UAV swarm framework is its ability to scale effectively as the number of agents increases. To evaluate the scalability of our proposed method, we conducted additional experiments with varying swarm sizes, specifically for $N\in\{10,20,30,40,50\}$ UAVs. Specifically, we focused on the greedy strategy, and selected attack success rate as the key performance indicator, which measures the proportion of time the attacker successfully disrupts a target UAV or link.
The results are presented in Fig. 7, where each subfigure compares the five defense methods for a fixed swarm size. Two general phenomena are evident across all subplots. First, as the swarm size increases, the attack success rate for all methods tends to rise. This is an expected outcome, as larger swarms present more complex coordination challenges and a larger attack surface for the adversary. Second, the success rate of greedy link attacks is consistently slightly higher than that of greedy node attacks across most methods and scales, which aligns with our earlier findings (e.g., in Table II) and suggests that disrupting communication links has a more cascading effect on the swarm’s integrity.
Despite these challenges, the key finding is that our proposed PG-FMADRL method demonstrates significantly superior scalability when compared to SOTA methods. In every tested scenario, PG-FMADRL consistently achieves the lowest attack success rate for both attack types. For instance, in the most challenging scenario with $N=50$, the attack success rates for baseline methods soar to $\geqslant$ 84.30% for node attacks and $\geqslant$ 96.58% for link attacks. In contrast, our method successfully constrains these rates to 31.00% and 39.96%, respectively. This robust scalability is attributed to its federated learning architecture. While the performance of independent learning methods degrades more sharply as each agent grapples with an increasingly complex local problem, our reward-weighted aggregation mechanism ensures that effective defense policies are efficiently shared and adapted across the entire, growing network. This allows the swarm to maintain a robust, collective defense posture even at large scales, confirming that PG-FMADRL is highly scalable for deployment in large-scale low-altitude networks.
(a) NN = 10
(b) NN = 20
(c) NN = 30
(d) NN = 40
(e) NN = 50
Figure 7: The attack success rate for different defense methods under greedy node and link attacks. Each subplot corresponds to a specific swarm size (NN).
@@ VI Conclusion and Future Work
In this paper, we proposed a novel FMADRL-driven MTD framework to proactively and adaptively mitigate DoS attacks in UAV swarm networks. By designing three lightweight and coordinated MTD mechanisms (e.g., leader switching, route mutation, and frequency hopping) and formulating the defense problem as a multi-agent POMDP, our approach enables each UAV to autonomously select defense actions based on local observations while benefiting from collaborative learning through reward-weighted federated aggregation. Extensive simulation results demonstrated that the proposed PG-FMADRL method significantly outperforms state-of-the-art baselines in terms of attack mitigation rate, recovery time, energy consumption, and defense cost, while maintaining robust mission continuity under various attack scenarios. By enhancing the security and robustness of UAV swarms, our work provides a foundational step toward the reliable and scalable deployment of low-altitude networks, supporting the vision of intelligent, resilient, and trustworthy low-altitude economic infrastructures.
Despite these promising results, our framework has limitations that open avenues for future research. A key limitation is that while the stochastic nature of our PG-FMADRL method inherently disrupts these adversaries, its resilience against a co-evolving, DRL-based adversary remains an open question. Another limitation is the reliance on a central aggregator which introduces a potential single point of failure. Therefore, future work should focus on enhancing our framework in two key directions. First, by evaluating it within a true adversarial learning context, potentially by modeling the interaction as a zero-sum game to develop more robust defense strategies. Second, by exploring more resilient aggregation models, such as decentralized or hierarchical federated learning, to eliminate the single point of failure.